OpenStudy (anonymous):

You save $15,000.00. You place one-third in a savings account earning a 4.6% APR compounded annually. You then invest one quarter of the remaining balance in a 3-year U.S. Treasury bond earning a 5.2% APR compounded annually and the rest in a stock plan. Your stock plan increases in value 3% the first year, decreases 8% in value the second year, and increases 6% in value the third year. What are the balances for each account by the end of the third year and the total gain on your original saved amount?

OpenStudy (btaylor):

For the 4.6% APR: you are investing 5000. the equation is 5000 x (1.046)^3 = 5722.23 For the Treasury bond: you are investing 2500. the equation is 2500 x (1.052)^3 = 2910.63 For the stocks: you are investing 7500. After 1 year, you have (1.03 x 7500) = 7725 After 2nd year, you have (7725 x 0.92) = 7107 After the third year, you have (7107 x 1.06) = 7533.42 So you gained 722.23 + 410.63 + 283.42 = $1416.28